zhockey55:

Identify the expression as an equation containing fractions, a fractional equation, or a proportion. x/12=17/3-2x/9-5+3x/12

jhonyy9:

what is the difference between ? 1/x + x^2 + x/3 = 7 and \[\frac{ 3x +2 }{ x +1 } = 7\] and (1/3 ) = (4/12) hope helped